Circular No. 8233

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003jf IN MCG +11-22-9 Further to IAUC 8226, T. Boles reports the discovery, on unfiltered

CCD images taken on Oct. 26.785 and 27.750 UT, of an apparent supernova (mag 18.4) located at α = 17^{h}53^{m}33^{s}.45, δ = +64^{o}10′12′′.1 (equinox 2000.0), which is ≈ 4′′.0 east and 11′′.1 south of the center of MCG +11-22-9. SN 2003jf is not present on his images from Sept. 13, 25, and Oct. 16 (limiting mag 19.5) or on Palomar Sky Survey red or blue plates.

SUPERNOVA 2003jc IN MCG –01-58-18 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003jc (cf. IAUC 8232), obtained by M. Calkins on Oct. 26.22 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows a blue, almost featureless continuum. There are nascent P-Cyg lines of hydrogen and helium, implying that this is an early type-II supernova.

1937 UB (HERMES) P. Pravec and P. Kušnirák, Ondřejov Observatory; B. Warner, Col-

orado Springs, CO, U.S.A.; R. Behrend, Geneva Observatory; A. W. Harris, Space Science Institute; A. Oksanen, Jyväskylä, Finland; D. Higgins, Canberra, Australia; R. Roy, Blauvac, France; C. Rinner, Ottmarsheim, France; C. Demeautis, Village-Neuf, France; F. van den Abbeel, Vesqueville, Belgium; A. Klotz, Guitalens, France; N. Waelchli, Observatoire François-Xavier Bagnoud, St-Luc, Switzerland; T. Alderweireldt, ’s-Gravenwezel, Belgium; V. Cotrez, St-Hélene, France; and L. Brunetto, Pitalugue, France, report that their photometric observations obtained during Oct. 16–26 show a lightcurve period for Hermes of 13.892 ± 0.006 hr with amplitude 0.07 ± 0.01 mag. A smooth shape of the lightcurve with only a single confirmed periodicity suggests that it is most likely the combined rotational lightcurve of two bodies synchronously locked to the orbital period (cf. IAUC 8227). The low amplitude implies that the two components seen by radar are physically separate orbiting bodies. No attenuations caused by mutual occultation/eclipse events were detected, implying that the mutual orbital plane was inclined to both the line of sight and the direction to the sun, or much less likely, that the orbital period is different from the detected rotation period. Continued monitoring for mutual occultations, as the viewing geometry changes, is warranted. The system’s total absolute magnitude is H = 17.55 ± 0.10.

2003 October 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green